# BSTcontrol

### v2

@@ 6.3.3 Joint Scheduling of Horizontal and Vertical Scalability
The integration of horizontal expansion and vertical enhancement offers a promising method to improve computational capability. Horizontal expansion allocates tasks across multiple devices or nodes while vertical enhancement increases the computational power of individual units. A joint scheduling framework coordinates resource allocation across both dimensions to achieve balanced workload distribution and reduced latency, which is essential for the effective deployment of large AI models in dynamic wireless environments [302, 303].
@@ 6.4 Security and Privacy Issues
While large AI models offer promising perspectives for the next-generation wireless networks, this integration presents significant security and privacy challenges. As these models handle large amounts of data and sensitive information, ensuring robust security and privacy protection mechanisms is crucial [279]. In the following parts, we discuss the security and privacy issues, as well as the defenses.
@@ 6.4.1 Security Issues
Large AI models are vulnerable to various security attacks. One well-known attack is the data poisoning attack, which involves injecting malicious data into the training dataset, causing the models to produce unreasonable and harmful results [280]. Another type of attack is the backdoor attack, which aims to embed a backdoor into the large AI models. When triggered with specific inputs, these backdoors can cause the AI models to perform unethical and illegal actions [281]. Moreover, the explainability of large AI models is crucial for the security of wireless communications. A lack of explainability makes it difficult to understand the underlying operation principles and inference processes, potentially leading to anomalies in communications that are challenging to audit and debug. As wireless communications infrastructure is a foundational building block of society, unexpected and harmful outputs could result in significant and unbearable losses at both social and economic levels. Therefore, it is essential to enhance the security of large AI models. Designing more explainable structures is a promising step towards achieving this goal. Additionally, ensuring transparency and security in the generation, collection, and cleaning of training data is crucial.
@@ 6.4.2 Privacy Issues
WLAM systems face significant privacy concerns. These concerns primarily arise from the vast amount of sensitive information these models process and the potential for unintended data exposure [282]. During the training and fine-tuning processes, large amounts of data from communication infrastructure and user equipment are utilized, which may contain vulnerabilities. Sensitive data can be at risk of leakage through methods like chain-of-thought prompting [283]. Additionally, commercial providers of large AI models often collect personal data and prompts fed to them, such as user location, user identity, and device information. The exposure of such private and sensitive data of numerous users poses threats from both service providers and attackers. As current large AI models lack standardized execution protocols and security constraints, it is crucial to develop corresponding data security principles and specifications for model design, implementation, and regulatory compliance. To ensure robust privacy protection in next-generation intelligent communications, establishing these standards and effectively enforcing security measures are essential.
@@ 6.5 Summary and Insights
In summary, the integration of large AI models into wireless communication offers substantial opportunities alongside significant challenges. Critical areas for progress include developing diverse, high-quality datasets, creating scalable and flexible network architectures, implementing energy-efficient AI strategies, and ensuring robust security and privacy measures. Addressing these issues will necessitate innovative solutions and interdisciplinary collaboration. Future research could focus on designing adaptive AI models capable of efficiently operating in dynamic environments while safeguarding data security and privacy. Overcoming these high-level challenges will unlock advanced AI-driven communication systems and a new era of wireless innovation.
@@ 7 Conclusions
A comprehensive exploration of WLAM for 6G and beyond has been presented in this survey, encompassing fundamentals, applications, challenges, and future directions. The synergistic potential of WLAM and wireless communication has been emphasized, with a particular focus on their mutual enhancement. The analysis has covered core characteristics, key applications in network optimization and resource management, as well as the integration of emerging technologies. Critical challenges, such as issues related to data, architecture, energy, and security, have also been examined. It is highlighted that realizing the full potential of WLAM requires dedicated research efforts in areas such as model efficiency, distributed learning, and robust security techniques. Ultimately, WLAM is envisioned as transformative technology that could revolutionize 6G and future wireless communication systems. Intelligence, adaptability, and efficiency are expected to be key drivers for enhanced digital experiences and unprecedented connectivity. Continued innovation in this dynamic field is deemed essential in harnessing the full power of WLAM for shaping the AI-native future of 6G and beyond.

### v4

@@ 6.3.4 Joint Scheduling of Horizontal and Vertical Scalability
To address the thermal throttling constraints on single edge devices, the integration of horizontal and vertical scalability offers a promising solution. While vertically boosting the computational power of an edge node (e.g., via dynamic frequency scaling or NPU acceleration) improves single-task acceleration, horizontal expansion offers superior network-wide efficiency. By offloading computation across various devices by using techniques like splitting inference, this method successfully distributes both thermal load and power consumption. A joint scheduling framework coordinates this allocation to achieve balanced workload distribution. This ensures that no single device exceeds its thermal design power limit while maintaining low latency, which is essential for the sustainable deployment of large AI models in dynamic wireless environments [razavi2024talescalesreconcilinghorizontal, li2020elastic].
@@ 6.4 Security and Privacy Issues
While Sections 3.1.5 and 4.4 respectively detailed the dual relationship of using AI for PLS and PLS for AI, this section provides a holistic overview of the broader security and privacy challenges confronting the entire WLAM paradigm. In the following parts, we discuss the security and privacy issues, as well as the defenses.
@@ 6.4.1 Security Issues
Large AI models are vulnerable to various security attacks [yao2024survey]. One well-known attack is the data poisoning attack, which involves injecting malicious data into the training dataset, causing the models to produce unreasonable and harmful results [wan2023poisoning]. Another type of attack is the backdoor attack, which aims to embed a backdoor into the large AI models. When triggered with specific inputs, these backdoors can cause the AI models to perform unethical and illegal actions [li2021hidden]. Moreover, the explainability of large AI models is crucial for the security of wireless communications. A lack of explainability makes it difficult to understand the underlying operation principles and inference processes, potentially leading to anomalies in communications that are challenging to audit and debug. As wireless communications infrastructure is a foundational building block of society, unexpected and harmful outputs could result in significant and unbearable losses at both social and economic levels. Therefore, ensuring the security of large AI models is paramount. Beyond patching specific vulnerabilities, future work needs to establish a comprehensive trustworthy AI framework for 6G, integrating formal verification methods to mathematically guarantee the safety and robustness of WLAM decisions in critical infrastructure.
@@ 6.4.2 Privacy Issues
WLAM systems face significant privacy concerns arising from their deployment architecture and interaction modes [brown2022does]. Unlike traditional on-device processing, the massive resource demands of WLAMs often necessitate cloud-based deployment, leading to data sovereignty risks where sensitive user context (e.g., location, channel logs) must be transmitted to third-party model providers. Furthermore, the inference process itself presents unique vulnerabilities; techniques like CoT prompting may inadvertently leak private intermediate reasoning steps or training data snippets in the output [xiang2024badchain]. Additionally, the lack of standardized privacy protocols for agentic AI creates loopholes where private network configurations could be accessed by unauthorized entities during API calls. To address these systemic risks, future research must establish privacy-preserving inference frameworks leveraging trusted execution environments [li2025teeslice] and homomorphic encryption [de2024privacy], ensuring that high-level intelligence can be delivered without compromising data confidentiality during the service process.
6.5 Channel Impact on Model Training
A foundational challenge that permeates the entire WLAM paradigm is the profound and multifaceted impact of the physical wireless channel on AI model training and operation. While many AI algorithms are developed assuming ideal data and reliable connections, the wireless medium is actually characterized by noise, fading, latency, mobility induced dynamics, and limited bandwidth which fundamentally alters their performance. This challenge is not confined to a single domain but is a cross-cutting issue affecting both the applications of large AI models for wireless systems (Section 3) and the wireless frameworks enabling distributed large AI models (Section 4).
6.5.1 Impact on Application-Specific Models
The challenges of the wireless channel manifest directly in the training process of large AI models designed for specific communication tasks.
The first challenge is the disruption of gradient-based training by the channel. In tasks like end-to-end transceiver design, the wireless channel acts as a stochastic, non-differentiable layer within the neural network. This physical reality makes standard gradient-based training processes fundamentally nonviable, as it breaks the chain of gradient back-propagation required for optimization [accessdecoding]. Consequently, the entire training paradigm must be re-designed, forcing the adoption of more complex strategies such as learning a differentiable surrogate channel model. These alternative strategies themselves introduce new challenges in terms of convergence stability and computational overhead [10104549].
Secondly, dynamic environments pose severe challenges to training convergence and generalization. For applications in mobile environments, such as channel prediction and beamforming, the non-stationary nature of the channel poses a severe threat to the training process. A model trained on channel data from one time-instance or location may fail to generalize moments later as the user moves. This dynamic nature means that a static, offline training approach based on simple data fitting is insufficient. It necessitates a fundamental shift in the training methodology itself. Instead of relying solely on massive datasets to implicitly learn the dynamics, more advanced paradigms are required that explicitly model the underlying physical principles of the mobile channel. For example, recent schemes leverage physics-inspired approaches such as ODE to create models whose internal structure is designed to represent and predict the continuous-time characteristics of the channel [10351043]. This transforms a straightforward supervised training task into a much more complex problem of training a physics-grounded dynamical system.
Lastly, the training process is often compromised by the train-test mismatch caused by imperfect CSI. Many AI models for tasks such as resource allocation are trained on datasets generated with the assumption of perfect CSI. However, in practice, CSI is always imperfect due to estimation errors, noise, and feedback delays. A model trained under such idealized assumptions may converge perfectly to a solution that is optimal for the simulation, but it will lead to poor generalization and performance degradation in in reality. Overcoming this issue requires models that are robust to these unseen perturbations. For instance, advanced training strategies like model-driven meta-learning can alleviate this problem by optimizing the search trajectory itself [zhurobustnetwork], enabling the model to find solutions that are inherently resilient to the impact of channel estimation errors.
To synthesize these solutions, future research must prioritize environment-aware adaptive training frameworks. By integrating differentiable channel modeling with physics-informed constraints directly into the optimization loop, WLAMs can bridge the gap between static model design and dynamic physical realities, ensuring generalization despite the stochastic and non-differentiable nature of the physical layer.
6.5.2 Impact on Distributed Training Frameworks
The impact of the channel is equally critical for the performance and stability of the distributed training frameworks.
Firstly, the channel introduces communication bottlenecks that directly impede the training process. Distributed frameworks like FL and SL require frequent exchange of large data payloads such as model parameters, gradients, and activations. The limited bandwidth of wireless channels creates a communication bottleneck, directly increasing the time required for each training round. This is explicitly captured in FL delay models, where transmission time is inversely proportional to bandwidth [yang2020delay]. For SL, the cumulative nature of wireless latency over its sequential communication process creates an even more severe bottleneck for the overall training time.
Secondly, channel impairments introduce errors that corrupt the training process. A clear example is in AirComp, which strategically uses the channel for aggregation. Its accuracy is highly vulnerable to channel impairments, as both channel noise and fading distort the superposition of signals at the receiver, introducing errors into the computed result [cao2021optimized]. This physical layer imperfection directly translates into noise injected into the model training process, which can slow down convergence or increase the optimality gap of the final model.
Lastly, the wireless channel creates vulnerabilities to adversarial attacks that compromise the training process. A malicious actor can exploit the open nature of the wireless medium to launch attacks like jamming. A jamming signal can corrupt the transmission of gradients or model updates, effectively blinding the central server. Research has established a direct mathematical link between the communication MSE over a jammed channel and the training loss of a large language model, formally demonstrating that an attacker can compromise the integrity of the entire training process by manipulating the physical channel [federatedPHY].
Therefore, the future research pathway lies in establishing robust communication-learning integration. This requires designing aggregation protocols that are not only communication-efficient to bypass bandwidth bottlenecks but also mathematically resilient to physical layer impairments and adversarial jamming, thereby preserving the integrity of the training process against wireless instability.
6.6 Fundamental Trade-offs in WLAM
The integration of large AI models into wireless communication is not a simple matter of maximizing performance, but rather one of navigating fundamental and often conflicting trade-offs. These inherent tensions define the core challenges and future research landscape for WLAM. The following analysis explores three of the most critical trade-offs: the balance between performance and efficiency, the conflict between data utility and privacy, and the tension between autonomy and reliability.
6.6.1 Performance versus Efficiency
A primary trade-off exists between the exceptional performance promised by WLAMs and the immense computational and energy demands they impose. While large models can achieve unprecedented accuracy in tasks like channel prediction or semantic understanding, their operational cost challenges deployment on resource-constrained wireless edge devices. Consequently, solutions like model quantization, pruning, or the joint scheduling of horizontal and vertical scaling are proposed. But they represent an inherent compromise, where peak model capability is deliberately traded for the energy efficiency and operational feasibility required in real-world wireless networks. This fundamental conflict indicates that a central future research direction for WLAM should be the pursuit of novel algorithms and hardware-software co-designs that push the Pareto frontier of this trade-off, enabling greater performance with less efficiency cost.
6.6.2 Data Utility versus Privacy
The second fundamental conflict arises between the need for high-fidelity data to ensure model utility and the necessity to protect user privacy. While traditional privacy research like FL focuses on decentralized training without raw data sharing, WLAMs introduce severe intrinsic privacy risks stemming from the models themselves. Specifically, the vast parameter space required for high performance creates a memorization-utility trade-off. The model’s capacity to learn complex wireless patterns also enables it to unintentionally memorize sensitive training samples, such as user mobility traces or unique channel fingerprints. This leaves the system vulnerable to model inversion or membership inference attacks, where adversaries can reconstruct private training data solely by querying the deployed model. Furthermore, a new issue emerges during the inference phase of agentic AI. To achieve high utility, agents require rich contextual prompts, which inherently exposes user intent and network state to the model provider. This exposure is compounded by prompt injection vulnerabilities, where malicious actors exploit the model’s flexible instruction following ability to bypass safety filters and attain confidential network configurations. Therefore, future research should focus on model-centric defenses such as machine unlearning [li2025machine], inference-time privacy guardrails [tong2025inferdpt], and robust alignment [kumar2025robustness] to balance these conflicting demands.
6.6.3 Autonomy versus Reliability
Finally, a critical trade-off emerges between the drive for network autonomy through agentic AI and the need for absolute reliability in critical infrastructure. The vision of a WLAM acting as an autonomous agent to manage a 6G network must be balanced against the severe risks posed by the flaws of large AI models. The potential for a WLAM to generate hallucinated network configurations, misinterpret high-level intents, or fall prey to backdoor attacks represents a profound threat to the stability and security of communication systems. Ensuring the reliability of these powerful but inherently probabilistic models is therefore a paramount challenge. Future research must prioritize the development of inherently explainable architectures to enable auditing, alongside trustworthy agentic frameworks that integrate formal verification and human-in-the-loop mechanisms to guarantee safe operation in critical infrastructure.
@@ 6.7 Summary and Insights
In summary, this section has systematically analyzed the primary obstacles confronting WLAMs, progressing from specific technical domains to fundamental trade-offs. The challenges in creating datasets, designing network architectures, ensuring efficiency, and safeguarding security are not isolated issues but reflect deeper conflicts between performance and efficiency, data utility and privacy, and autonomy and reliability. Navigating these tensions requires a paradigm shift from isolated technical fixes to holistic, interdisciplinary solutions. Future research must focus on coordinating these conflicting demands by integrating hardware-software co-design, trustworthy agentic frameworks, and intrinsic privacy defenses into a unified architecture. The future of WLAM depends on developing these holistic solutions to bridge the gap between theoretical capabilities and operational realities.
@@ 7 Conclusions
This survey has provided a comprehensive exploration of the WLAM, charting its fundamentals, applications, and the path toward an AI-native 6G. Our key finding is the deeply synergistic and dual relationship where large AI models are revolutionizing wireless systems, and advanced wireless technologies are, in turn, enabling the deployment of large AI models. We have demonstrated that this transformation extends across the entire communication stack. This includes enhancing the physical layer with intelligent transceiver design, optimizing the network layer through automated resource management, and redefining efficiency at the semantic layer, ultimately creating a new agentic layer for autonomous network operations. Furthermore, we identified that future breakthroughs will be driven by the integration of emerging technologies, such as PINNs and next-generation sequence models, which promise to enhance data efficiency and computational performance.
While the potential is immense, significant challenges remain. To guide future work, we highlight several critical research directions. First is the pressing need for standardized multi-modal wireless datasets, which are essential for benchmarking and developing robust models that can fuse diverse data sources like CSI and sensor data. Second, a concerted effort is required to develop edge-native AI, focusing on novel, highly efficient model architectures and hardware-software co-design to overcome the severe resource constraints of wireless devices. Third, as networks become more autonomous, research must prioritize trustworthy autonomy, by developing methods to enhance the reliability, explainability, and security of WLAMs against threats like model hallucinations and backdoor attacks. Finally, the deeper fusion of physics-informed AI with large models presents a fertile ground for creating more generalizable and data-efficient solutions grounded in the fundamental principles of wireless communication. Continued innovation in these areas is essential to harness the full power of WLAMs, shaping the intelligent, adaptive, and efficient future of 6G and beyond.
While the potential is immense, significant challenges remain. To guide future work, we highlight several critical research directions. First, there is a need for standardized multi-modal wireless datasets and synthetic data generation methods, which are essential for benchmarking and developing robust models that can fuse diverse data sources like CSI and sensor data. Second, a concerted effort is required to develop edge-native AI, focusing on novel, highly efficient model architectures and hardware-software co-design to break the “memory wall” and overcome the severe resource constraints of wireless devices. Third, as networks become more autonomous, research must prioritize trustworthy autonomy and formal verification, by developing methods to enhance the reliability, explainability, and security of WLAMs against threats like model hallucinations and backdoor attacks. Fourth, addressing intrinsic privacy risks requires exploring model-centric defenses, such as machine unlearning, to prevent the unintended memorization of sensitive user data. Finally, the deeper fusion of physics-informed AI with large AI models presents a fertile ground for creating more generalizable and data-efficient solutions grounded in the fundamental principles of wireless communication. Continued innovation in these areas is essential to harness the full power of WLAMs, shaping the intelligent, adaptive, and efficient future of 6G and beyond.